OpenStudy (anonymous):

The Sum of the Series : \[\frac{ 1 }{ 2 }-\frac{ 3 }{ 2^2 }+\frac{ 1 }{ 2^3}-\frac{ 3 }{ 2^4 }+......+\infty \]

OpenStudy (anonymous):

@satellite73 @mahmit2012 @asnaseer @eigenschmeigen

OpenStudy (ash2326):

Split the series \[(\frac 1 2 +\frac 1 {2^3}+....)-3(\frac 1 {2^2}+ \frac 1 {2^4}...)\] Now these both are infinite Geometric series. Could you find the sum of each ?

OpenStudy (asnaseer):

if you factor out another 1/2 from the negative series you should spot something

OpenStudy (anonymous):

break in to two parts

OpenStudy (anonymous):

oh, what everyone else said, sorry

OpenStudy (anonymous):

(2-3)*geo series